Grade 12th passPhysical Chemistry

Calculate the ph of NaoH​....................!!!!!!!!!1

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Pawan Prajapati
5 Years ago
Answer:

Alternatively we can calculate pOH = - log [OH-]= - log [10-1] = 1 and remembering that pH + pOH = 14 we get pH = 14 - 1 = 13. follow me

Thanks